OpenStudy (anonymous):

Solve 2x2 - 8x = -7.

OpenStudy (solomonzelman):

\(\large\color{blue}{ 2x^2 - 8x = -7 }\) \(\large\color{blue}{ 2(x^2 - 4x) = -7 }\) \(\large\color{blue}{ 2(x^2 - 4x+4-4) = -7 }\) \(\large\color{blue}{ 2(x^2 - 4x+4) -8= -7 }\) \(\large\color{blue}{ 2(x^2 - 4x+4)= 1 }\) \(\large\color{blue}{ (x^2 - 4x+4)= 1/2 }\) \(\large\color{blue}{ (x-2)^2= 1/2 }\) \(\large\color{blue}{ x-2= ±\sqrt{1/2} }\)

OpenStudy (solomonzelman):

Oops, can you please refresh, I clicked the plus minus symbols and it shows question marks instead. Sorry about this inconvenience.

OpenStudy (anonymous):

Wrong thing

OpenStudy (anonymous):

Hold on

OpenStudy (loser66):

lol

OpenStudy (anonymous):

My answer choices

OpenStudy (solomonzelman):

this is 12th grade (I would try to withhold myself from putting a question and an exclamation mark right after, at the end)